Use Lagrange’s mean value theorem to prove that


x-x^2/2《log(1+x)《x-x^2/2(1+x) if x》0.
1
Expert's answer
2020-03-02T16:33:54-0500

"x-\frac{x^2}{2}\leq\log (1+x)\leq x-\frac{x^2}{2(1+x)}, \ \ \quad \forall \ x\geq0"

Proof:

Lagrange's Mean Value Theorem:

Let "f: [a,b]\rightarrow \mathbb{R}" be a continuous function, differentiable on the open interval "(a,b)"  . Then there exists some  "c\in(a,b) \" such that "f^\prime(c)=\frac{f(b)-f(a)}{b-a}" .



"f(x)=\log(1+x)+\frac{x^2}{2}"

This function is continuous in "[0,b]" , differentiable in "(0,b)," for all "b\geq0."

Using Lagrange’s Mean Value Theorem for function "f(x)", we have:

"\exists c\in (0,b): \ \frac{f(b)-f(0)}{b-0}=f^\prime(c)"

"\frac{f(b)-f(0)}{b-0}= \frac{(\log(1+b)+\frac{b^2}{2})-(\log(1)+0)}{b}= \frac{\log(1+b)+\frac{b^2}{2}}{b}=f^\prime(c)=\frac{1}{1+c}+c"

"\frac{\log(1+b)+\frac{b^2}{2}}{b}=\frac{1}{1+c}+c=1+\frac{c^2}{1+c}\geq1 \Rightarrow \frac{\log(1+b)+\frac{b^2}{2}}{b}\geq1 \Rightarrow \log(1+b)+\frac{b^2}{2}\geq b"

"\Rightarrow \log(1+b)\geq b-\frac{b^2}{2}, \forall b\geq0"


So, "x-\frac{x^2}{2}\leq\log (1+x), \ \ \forall \ x\geq 0"


"f(x)=\log(1+x) +\frac{x^2}{2(1+x)}"

This function is continuous in "[0,b]" , differentiable in "(a,b)" for all "b\geq 0."

Using Lagrange’s Mean Value Theorem for function "f(x)" , we have:

"\exists c\in (0,b): \ \frac{f(b)-f(0)}{b-0}=f^\prime(c)"

"\frac{f(b)-f(0)}{b-0}= \frac{(\log(1+b)+\frac{b^2}{2(1+b)})-(\log(1)+0)}{b}= \frac{\log(1+b)+\frac{b^2}{2(1+b)}}{b}=f^\prime(c)="

"=\frac{1}{1+c}+\frac{2c\times 2(1+c)-2c^2}{4(1+c)^2}=\frac{2c^2+8c+4}{4c^2+8c+4}=1-\frac{2c^2}{4c^2+8c+4} \leq 1"

"\frac{\log(1+b)+\frac{b^2}{2(1+b)}}{b}\leq 1\Rightarrow {\log(1+b)+\frac{b^2}{2(1+b)}}\leq b 
\Rightarrow {\log(1+b)\leq b- \frac{b^2}{2(1+b)}}, \ \forall b\geq 0"


So, "\log (1+x)\leq x-\frac{x^2}{2(1+x)}, \ \ \quad \forall \ x\geq0"